Using the following data on GDP per capita in the United Kingdom, compute the average annual growth rates during each of the 5 decades and the average annual growth rates over the 50 years

3. 6 points Using the following data on GDP per capita in the United Kingdom, compute the average annual growth rates during each of the 5 decades and the average annual growth rates over th 50 vears Year Real GDP per capita 1960 1970 1980 1990 2000 2010 (2011 dollars) $11,879 $15,308 $18,571 $24,799 $31,640 $35,689
Answer #1

(A) Average annual growth rate in each decade is computed as follows.

1960-1970 = (15,308 / 11,879) - 1 = 1.2887 - 1 = 0.2887 = 28.87%

1970-1980 = (18,571 / 15,308) - 1 = 1.2132 - 1 = 0.2132 = 21.32%

1980-1990 = (24,799 / 18,571) - 1 = 1.3354 - 1 = 0.3354 = 33.54%

1990-2000 = (31,640 / 24,799) - 1 = 1.2759 - 1 = 0.2759 = 27.49%

2000-2010 = (35,689 / 31,640) - 1 = 1.1280 - 1 = 0.1280 = 12.80%

(B) Average annual growth rate during 50 years is computed as follows.

1960-2010 = (35,689 / 11,879) - 1 = 3.0044 - 1 = 2.0044 = 200.44%